# Dark matter: red or blue?

### v2

@@ 3 Gravitational Interactions
$\begin{split}\frac{d\sigma_{\parallel\parallel}}{d\Omega}=\frac{\lambda^{4}(K+P)^{4}}{16(2\pi)^{2}P^{2}(1-\text{cos}\theta)^{2}}\\ \frac{d\sigma_{\perp\perp}}{d\Omega}=\frac{\lambda^{4}(K+P)^{4}\text{cos}^{2}\theta}{16(2\pi)^{2}P^{2}(1-\text{cos}\theta)^{2}}\end{split}$ (10)
@@
Figure 5: The differential $\gamma\chi$ scattering cross section via gravitational interaction for the $M_{\chi}=10^{19}$ GeV as a function of photon scattering angle for two photon polarisations. (Note the $10^{247}$ scale)
@@
$\bullet$ Large angle scattering leads to non-trivial polarization dependence, Fig. 5. Since the cross-section is small, it is highly unlikely that we will be able to measure flux alteration due to $\gamma\chi$ gravitational scattering in the foreseeable future. However, polarisation effects may offer different (improved by $\sim$ 100 orders of magnitude) measurement sensitivities and systematics compared to flux measurements  [^5] Polarisation effects depends on amplitude linearly, while flux depends quadratically.

### v3

@@ 3 Gravitational Interactions
$\begin{split}\frac{d\sigma_{\parallel\parallel}}{d\Omega}=\frac{\lambda^{4}(K+P)^{4}\text{cos}^{2}\theta}{16(2\pi)^{2}P^{2}(1-\text{cos}\theta)^{2}}\\ \frac{d\sigma_{\perp\perp}}{d\Omega}=\frac{\lambda^{4}(K+P)^{4}}{16(2\pi)^{2}P^{2}(1-\text{cos}\theta)^{2}}\end{split}$ (10)
@@
Figure 5: The differential $\gamma\chi$ scattering cross section via gravitational interaction for the $M_{\chi}=10^{19}$ GeV as a function of photon scattering angle for two photon polarisations, using $E_{\gamma}=200$ GeV. (Note the $10^{27}$ scale)
@@
$\bullet$ Large angle scattering leads to non-trivial polarization dependence, Fig. 5. Since the cross-section is small, it is highly unlikely that we will be able to measure flux alteration due to $\gamma\chi$ gravitational scattering in the foreseeable future. However, polarisation effects may offer different (improved by $\sim$ 15 orders of magnitude) measurement sensitivities and systematics compared to flux measurements  [^5] Polarisation effects depends on amplitude linearly, while flux depends quadratically.